Tautitanna:

Marneshia walked three over eight of a mile in three over five of an hour. What equation can be used to calculate her unit rate in miles per hour? (5 points) three fifths over three eights = five eighths mile per hour three-eighths over three fifths = one and three-fifths miles per hour three fifths over three eights = one and three-fifths miles per hour three-eighths over three fifths = five-eighths mile per hour

Tautitanna:

help plz

AZ:

when we have something like `x` miles in `y` hours and we're just trying to get the UNIT RATE we have to divide x by y and that would tell us `x/y`miles in `1` hour and that would be the unit rate because now we can say He went x/y miles per hour

AZ:

so for your question we have 3/8 miles in 3/5 hours so how many miles would that be in 1 hour

AZ:

remember when you're dividing fractions: \(\dfrac{a}{b} \div \dfrac{c}{d} = \dfrac{a}{b} \times \dfrac{d}{c}\) \( = \dfrac{ad}{bc}\)

Tautitanna:

three-eighths over three-fifths = one and three-fifths miles per hour is this it.

AZ:

the first part was correct so let's look at the division again, okay? when we're dividing two fractions, we have to take the second fraction and FLIP it and then we can multiply it so we have \(\dfrac{3}{8} \div \dfrac{3}{5} \) then we get \(\dfrac{3}{8} \times \dfrac{5}{3}\) can you multiply it now?

AZ:

(hint: you'll notice that the 3's can cancel out)

Tautitanna:

three-eighths over three-fifths = five-eighths mile per hour??

AZ:

that's correct

Tautitanna:

ok thx

AZ:

yw